# The physical basis of teleological evolutions

### v4

@@ 5.5 Impact on philosophy
Of course, the current theory of teleological evolutions would rehabilitate an ancient notion of natural philosophy $\left[39\right]$. We should therefore be allowed to look for the possibility that there are other relations between this theory and philosophy. As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
@@
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
@@
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[13\right]$, and the present theory is a mutual causality view.
We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond to the causal and mutual causality views of evolutions, respectively.
For what concerns reductionism vs monism, we note that deriving the teleological character of the decohered evolutions of the living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view. It presupposes that what we see around and within us is essentially inseparable from the whole.
@@
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, are the causes of natural phenomena and constitute their unity” $\left[40\right]$. In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. The identity between the fundamental physical laws since the beginning of time and the ideas of our mind today is formally reminiscent of the identity (and mutual causality) between the number of the drawer with the ball (the problem-setting) randomly selected by Bob’s measurement at the beginning of the quantum algorithm and the number of the drawer with the ball (the solution of the problem) selected by Alice’s measurement at the end of it. It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
@@
Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, when it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
@@
The above should also go along with Fritjov Capra’s $\left[41\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
@@
(ii) We do not need to be anthropocentric. This mutual causality would also be between physical laws and spatial intelligence, often superior to our in non-human animals. The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality). Indeed a quantum speedup in the development of spatial intelligence would imply the mutual causality in question.
@@
(iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[42\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
@@
(iv) It is natural to wander whether the present time-symmetric quantum causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
@@
(v) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[43\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today? Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[44\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it? We would not feel like dismissing these analogies as sheer coincidences. As a matter of fact, in idealistic philosophy – in Schelling – aesthetic insight would reproduce the fundamental mechanisms of the production of life.
@@ 6 Research prospects
(i) Look for a synthesis between the present notion of the time-symmetric quantum causal loops and time-symmetric quantum mechanics $\left[20-24\right]$ to which it is obviously inspired.
(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (the mostly autonomous evolution of a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
(iii) The mechanism described in (ii) might apply to other life processes for which the presence of a quantum behavior is suspected, for example, to the possible quantum behavior of the brain. Although mostly autonomous from the rest of the universe, this behavior could derive from that of a quantum universe by tracing out everything external to the brain. Note the radical difference between this view and the usual one where people looks for an autonomous quantum behavior in the macroscopic, warm and wet brain. By the way, we could introduce a distinction between two kinds of quantum behaviors: autonomous, like those of microscopic systems sufficiently isolated from all the rest, and derived, namely quantum behaviors of even macroscopic subsystems obtained from that of a quantum universe – the system – by tracing out all the rest.
(iv) Look for the possible impact of the teleological view on quantum cosmology. According to current assumptions, even the inanimate part of the latter, always an evolution toward life, should be endowed with quantum speedup and thus have teleological character.
(v) The Anthropic Principle is limited to stating the correlation between the “value” of life and the value of the fundamental physical constants since the beginning of time. One could also look for a correlation between the value of life and the character of the fundamental physical laws since the beginning of time. Of course, also these would have an attractor in the goal of creating life.
@@
(II) Toy models of teleological Darwinian processes. These evolutions could be either approximated as pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
(III) An application of the time-symmetric quantum causal loops to quantum self-organization. An example is studying the teleological formation of a self-replicating system. There seem to be an enormous jump of complexity in the transition from a soup of life-precursor molecules into a self-replicating system. The idea that this transition has an attractor in the full fledged self-replicating system can therefore be attractive.
@@ 7 Conclusions
The work is, of course, highly unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
@@
We have (trivially) shown that the sheer existence of the optimal quantum speedup of Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
@@
Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic. Therefore, Grover algorithm must host the quantum superposition of the causal loops in question.
However, being causal in character as that of any other quantum process, its customary quantum description cannot describe causal loops; we ascribe this to its incompleteness.
@@
Before going to the way of remedying this incompleteness, we must eliminate another, trivial, one. The customary quantum description of quantum algorithms is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is incomplete since it lacks the initial measurement.
Thus, as a first step, we extend the customary quantum description of the process of solving the problem to the upstream process of setting the problem, by adding the initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings. We have thus a unitary evolution between two one-to-one correlated measurement outcomes that does not change the problem-setting selected by the initial measurement. This is a key feature of quantum algorithms, not visible in their customary quantum description777 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting. . We call this kind of process the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
The second step is to change the causal structure of the ordinary quantum description of the quantum algorithm by time-symmetrizing the latter. We time-symmetrically share between the initial and final measurements the selection of the information that specifies the sorted out pair of correlated measurement outcomes, in a quantum superposition of all the possible ways.
Each time-symmetrization instance, mathematically, yields the unitary evolution between the two measurement outcomes back again, but the structure of causality along it changes. It becomes a causal loop that changes its own past. Relativized with respect to the problem-solver (who should be concealed from any outcome of the initial measurement), the causal loop in question becomes one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops. Of course, the time-symmetrization leaves the ordinary quantum description mathematically unaltered but changes its single causal process into a quantum superposition of causal loops.
Until now, the time-symmetrization in question has been imposed by the necessity of reproducing, in the quantum description of Grover algorithm, the causal loops logically implied by the sheer existence of its (optimal) quantum speedup. However, as we have seen, the same time-symmetrization is also redundantly imposed by a symmetry implicit in any unitary evolution between an initial and a final, one-to-one correlated, measurement outcomes that does not change the eigenvalue selected by the initial measurement. As a consequence, any such evolution is a quantum superposition of time-symmetric causal loops.
If the unitary evolution builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement, like in the case of Grover algorithm, in each time-symmetrization instance the dimension of the Hilbert space in which the search for one-to-one correlation takes place reduces. This creates the room for a quantum speedup. There is a quantum speedup even in the case that the search for one-to-one correlation is random.
The process is a quantum superposition of time-symmetric causal loops also in the case that the two observables are already maximally entangled at the beginning of it. It is the case of quantum nonlocality, where the quantum process starts with the spatial separation between two subsystems whose respective observables are already maximally entangled. Here the causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell either observable the state to assume in view of its future measurement. They play the role of the hidden variables envisaged by Einstein and the others to locally explain quantum nonlocality.
@@
The quantum causal loop description of quantum algorithms highlights their teleological character, namely their being evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again). This naturally raises the question whether the time-symmetric quantum causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
@@
This immediately raises the problem of explaining a quantum speedup in evolutions of the living that can be the size of the Earth and have durations of billions of years. The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the only possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time we should adopt quantum cosmology, namely the assumption that the universe evolves in a unitary quantum way.
@@
We should now call into play the “Fine-tuned Universe” version of the Anthropic Principle. It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”. We assume for simplicity that this correlation is one-to-one.
@@
With the evolution of the universe quantum, also the correlation asserted by the Anthropic Principle becomes quantum. We are thus dealing with a unitary evolution that one-to-one correlates two initially unrelated observables (the value of the fundamental constants and the value of life observables) without changing the eigenvalue of the observable selected by the initial measurement. We have seen that this kind of unitary evolution can occur with quantum speedup and does so even in the case that the search for one-to-one correlation is random.
@@
The evolution of the universe toward the value of life should therefore take place with quantum speedup and thus be teleological in character. In the universe we are in, this would apply to the evolution of the universe toward sentient life.
@@
We have also seen that the evolution of life on Earth, in principle obtainable from that of the universe toward sentient life by tracing out all the rest, should inherit the teleological character of the latter evolution. In fact: (i) evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe and (ii) the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
In conclusion, the notion of the time-symmetric quantum causal loops turns out to be the common physical basis of such diverse physical phenomena as the quantum computational speedup, the consequent teleological character of quantum algorithms, quantum nonlocality, and – under a quantum cosmological interpretation of the Anthropic Principle – the teleological character of the evolutions of the living.
Of course, that of teleological evolution is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
Regarding the quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
Regarding the teleological character of the macroscopic evolutions of life, the situation becomes even more puzzling. We limit ourselves to noticing that this is a necessary consequence of the quantum cosmological assumption. Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience some of the strange characters of the microscopic quantum world.

### v6

@@ 4.5 Impact on philosophy
Of course, the current theory of teleological natural evolutions has a precise conceptual precedent in a very ancient notion of natural philosophy $\left[41\right]$. We should therefore be allowed to look for other possible relations between this theory and philosophy. As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
@@
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view777 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
@@
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$ is a mutual causality view.
We find a similar situation in the opposition between the mechanistic and teleological views of evolutions, which obviously respond, respectively, to the causal and mutual causality views of evolutions.
Regarding reductionism vs. monism, we note that deriving the teleological character of the decohered evolutions of living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view. It presupposes that the description of what we see around and within us is essentially inseparable from the description of the whole.
@@
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[42\right]$. In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. The identity between the fundamental physical laws since the beginning of time and the ideas of our mind today is formally reminiscent of the identity (and mutual causality) between the number of the drawer with the ball (the problem-setting) randomly selected by Bob’s measurement at the beginning of the quantum algorithm and the number of the drawer with the ball (the solution of the problem) selected by Alice’s measurement at the end of it. It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
@@
Also Plato’s notion that ideas – physical laws – are objective perfections could have a correspondence with the fine tuned universe version of the Anthropic Principle, the fact it asserts that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
@@
The above should also go along with Fritjov Capra’s $\left[43\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
@@
(ii) We do not need to be anthropocentric. This mutual causality would also be between physical laws and spatial intelligence, often superior to our in non-human animals. The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality). Indeed a quantum speedup in the development of spatial intelligence – and the consequent teleological character of this development – would imply the mutual causality in question.
@@
(iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[44\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
@@
(iv) It is natural to wander whether the present time-symmetric causal loops view of the universe’s evolution toward sentient life, where the sentient life stage is an attractor of the evolution of the universe toward it, could be related to Nietzsche’s intuition of an eternal return consequent to our will for it.
@@
(v) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[45\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today? Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[46\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it? We would not feel like dismissing these analogies as sheer coincidences. As a matter of fact, in idealistic philosophy – in Schelling – aesthetic insight would reproduce the fundamental mechanisms of the production of life.
Overall, it would seem that the current ”scientific” theory of teleological evolutions reevaluates the idealistic side of philosophy against the materialistic side. However, the latter would also have its revenge, since a scientific justification of idealism could be ascribed to a materialistic view.
To sum up, it seems clear that the contrast between the causal and mutual causality views touches a deep mechanism in the history of thought. In an inevitably biased way, we would ascribe it to the unifying power of the current theory of teleological evolutions.
@@ 5 Research prospects
(i) Look for a synthesis between the present notion of the time-symmetric causal loops and time-symmetric quantum mechanics $\left[19-23\right]$ to which it is obviously inspired.
(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem, although decohered, inherits the quantum speedup and consequent teleological character of the evolution of a purely quantum system. Of course, here, the quantum system stands for the universe and the subsystem for the living Earth subsystem.
Note that the present notion for which the decohered behavior of a subsystem inherits quantum features from the behavior of the quantum system might apply to other cases. Possible examples are the life processes suspected to have higher than classical efficiency and consciousness suspected to need quantum nonlocality, the quantum system being here, of course, the universe.
(iii) Look for a possible impact of the present teleological view on quantum cosmology. According to current assumptions, even the “inanimate” part of the universe’s evolution, an evolution toward life too, should be endowed with quantum speedup and thus have teleological character.
(iv) The Anthropic Principle is limited to asserting the correlation between the “value” of life today and the value of the fundamental physical constants since the beginning of time. One could also look for a relation between the value of life today and the character of the fundamental physical laws since the beginning of time. Of course, in the universe we are in, also the latter character should have an attractor in the goal of creating life.
@@
(II) Toy models of teleological Darwinian processes. These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
(III) An application of the time-symmetric causal loops to quantum self-organization. An example could be studying the teleological formation of a self-replicating system. There might be an enormous jump of complexity in the transition from a soup of life-precursor molecules to a self-replicating system. The idea that this transition has an attractor in the full fledged self-replicating system can therefore be attractive.
@@ 6 Conclusions
It might be useful to summarize the key points of our unconventional argument.
@@
We have trivially shown that the sheer existence of the quantum speedup of the optimal Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
On the one hand, being trivial, this claim is perfectly verifiable. On the other hand, a trivial demonstration of the existence of causal loops, a notion strongly questioned, should therefore be a given. This can be considered the fulcrum of the present work, and the source of its unconventionality. If, so simple, it has never been noticed outside the present retrocausal approach to the quantum speedup, it must be because a tacit no-retrocausality precept has prevented looking that way.
The following is the consequent line of thinking.
@@
Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic of the logical implication in question. Therefore, Grover algorithm must host the quantum superposition of the causal loops we are dealing with. However, of course, this cannot be in its quantum description that, being causal in character as that of any quantum process, cannot describe causal loops.
@@
It should be noted that the usual quantum description of quantum algorithms presents another problem. It is limited to the process of solving the problem where the quantum speedup mathematically resides and is clearly incomplete. In fact, it consists of the unitary computation of the solution and the measurement of the latter. It lacks the initial measurement.
Remedying this incompleteness is another key point of the work. Extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in the quantum superposition (or mixture) of all the possible problem-settings makes the quantum description of the quantum algorithm time-symmetric. The quantum description of Grover algorithm becomes that of a unitary evolution between an initial and a final one-to-one correlated measurement outcomes. Of course, being still causal, this extended quantum description cannot describe causal loops. However, time-symmetrizing it generates the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes. The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement. The selection performed by the final measurement should propagate backward in time until replacing the previous outcome of the initial measurement. This is a causal loops that changes its own past – the initial measurement outcome. It is indeed one of the causal loops logically implied by the sheer existence of the quantum speedup.
Of course, we should take the quantum superposition of all the possible time-symmetrization instances, each for each possible way of time-symmetrically sharing the selection of the pair of measurement outcomes between the initial and final measurements. This leaves the quantum description of the unitary evolution between the initial and final measurement outcomes mathematically unaltered but changes the description of the structure of causality along that evolution. The single causal process of the ordinary quantum description is replaced by the quantum superposition of causal loops logically implied by the sheer existence of the quantum speedup. Being generated by a time-symmetrization, we call these causal loops time-symmetric.
Note that the time-symmetrization in question turns out to be the only possible way of obtaining a quantum description of the quantum algorithm that describes the causal loops the quantum algorithm must host. It is therefore mandatory.
We have shown that this immediately generalizes to any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (the problem-setting in quantum algorithms). The ordinary quantum description of any such process is incomplete and is completed by time-symmetrizing it. This leaves the ordinary quantum description of the unitary part of the quantum process mathematically unaltered but replaces its single causal process by a quantum superposition of time-symmetric causal loops.
In particular, the generalization covers the case of quantum nonlocality, where the unitary evolution is the process of spatial separation of the two subsystems in a state of maximal entanglement between the respective observables. There is one-to-one correlation between the outcomes of measuring these observables in the same basis. Here, the time-symmetric quantum causal loops play the role of the hidden variables envisaged by Einstein and the others to complete the ordinary quantum description of quantum nonlocality.
@@
Eventually, the quantum causal loop description of quantum algorithms highlights their teleological character, that is, their being evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again). This naturally raises the question whether the time-symmetric causal loops we are dealing with could also be the physical basis of the teleological character of the evolutions of the living for which the notion of teleological evolution was conceived.
We believe that investigating this possibility is incumbent on the part of physics toward the natural sciences, after its having denied for centuries the possibility that teleological evolutions exist.
@@
This immediately raises the problem of explaining the presence of quantum features (the quantum speedup) in evolutions of the living that can be the size of the Earth and have durations of billions of years. The obvious impossibility that the quantum features of the microscopic world scale up to such levels suggests the only possible solution: that they scale down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time we should assume, with quantum cosmology, that the evolution of the universe is unitary quantum. The fact that, once given the amount of resources involved, the quantum speedup is related to the time taken by the quantum process, under suitable conditions makes it inheritable by a subsystem (a natural evolution) of the quantum system (the universe).
@@
Surprisingly, the “Fine-tuned Universe” version of the Anthropic Principle puts all the above things together. It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”. We assume for simplicity that this correlation is one-to-one.
The assumption that the evolution of the universe is unitary quantum, of course forces us to adopt a quantum interpretation of the Anthropic Principle. It is now the unitary evolution of the universe that builds up a quantum correlation between the value of the physical constants since the beginning of time and the value of life today.
This is formally identical to a quantum search algorithm that unitarily builds up a quantum correlation between the value of the problem-setting since its beginning and the value of the solution and its end.
@@
We have seen that any unitary evolution that builds up one-to-one correlated between an initial and a final measurement outcome without changing the eigenvalue selected by the initial measurement can occur with quantum speedup and does even if the search for that correlation is random. The latter is of course a conservative assumption; that there has been that building up by the evolution of the universe is what the Anthropic Principle tells us.
@@
In conclusion, the unitary evolution of the universe toward the value of life occurs with quantum speedup and is thus teleological. In the universe we are in, the same, of course, holds for the evolution of the universe toward life.
@@
The evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out all the rest, should inherit the teleological character of the latter evolution. In fact: (i) evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe and (ii) the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
In conclusion, we have seen that the ordinary, causal, quantum description of a class of processes involving quantum correlation is incomplete, exactly in the sense argued by Einstein and the others in their EPR paper, and is completed by time-symmetrizing it. This explains such diverse quantum phenomena as the quantum speedup and the consequent teleological character of quantum algorithms, quantum nonlocality, and, under a quantum cosmological interpretation of the Anthropic Principle, the teleological character of the evolutions of life. It also recovers a pre-scientific view of physical reality that had waned with the advent of modern science and its classical-causal view of physical phenomena.